# ProcessPainter: Learn Painting Process from Sequence Data

### v1

@@ Abstract.
The painting process of artists is inherently stepwise and varies significantly among different painters and styles. Generating detailed, step-by-step painting processes is essential for art education and research, yet remains largely underexplored. Traditional stroke-based rendering methods break down images into sequences of brushstrokes, yet they fall short of replicating the authentic processes of artists, with limitations confined to basic brushstroke modifications. Text-to-image models utilizing diffusion processes generate images through iterative denoising, also diverge substantially from artists’ painting process. To address these challenges, we introduce ProcessPainter, a text-to-video model that is initially pre-trained on synthetic data and subsequently fine-tuned with a select set of artists’ painting sequences using the LoRA model. This approach successfully generates painting processes from text prompts for the first time. Furthermore, we introduce an Artwork Replication Network capable of accepting arbitrary-frame input, which facilitates the controlled generation of painting processes, decomposing images into painting sequences, and completing semi-finished artworks. This paper offers new perspectives and tools for advancing art education and image generation technology.
@@ 2.1. Stroke-based Rendering
The problem of teaching machines “how to paint” has been extensively studied in the context of stroke-based rendering (SBR), which focuses on the recreation of non-photorealistic imagery through appropriate positioning and selection of discrete elements such as paint strokes or stipples(Hertzmann, 2003). Early SBR approaches involved either a greedy, iterative search or required user interaction to address the decomposition problem (Haeberli, 1990; Litwinowicz, 1997; Hertzmann, 2022). In recent years, recurrent neural networks (RNNs) and reinforcement learning (RL) have been extensively employed to generate strokes sequentially (Ha and Eck, 2017; Zhou et al., 2018; Xie et al., 2013; Singh et al., 2022). Moreover, integrating adversarial training (Nakano, 2019) has emerged as a viable method for generating non-deterministic stroke sequences. adversarial updates of actors, critics, and discriminators. Stylized Neural Painting(Zou et al., 2021; Kotovenko et al., 2021) proposed stroke optimization strategy that iteratively searches optimal parameters for each stroke and and is possible to be optimized jointly with neural style transfer. Although the aforementioned methods achieve the painting process from abstract to specific, these methods still significantly differ from the creative processes of human artists. This is because the painting processes of different artists and subjects vary significantly, making it challenging for a single method to simulate all. Recently, the emergence of Stable Diffusion(Rombach et al., 2022) technology has significantly improved the effectiveness of generative images. Exploring the generation of painting processes based on diffusion models is a promising new field. To this end, we propose ProcessPainter, which fine-tunes diffusion models using data from artists’ painting processes to learn their true distributions.
@@ 2.2. Text2image Diffusion Model
Recent studies have demonstrated that diffusion models are capable of generating high-quality synthetic images, effectively balancing diversity and fidelity. Models based on diffusion models or their variants, such as those documented in (Nichol et al., 2021; Ramesh et al., 2022; Saharia et al., 2022), have successfully addressed the challenges associated with text-conditioned image synthesis. Stable Diffusion (Rombach et al., 2022), a model based on the Latent Diffusion Model (Rombach et al., 2022), incorporates text conditioning within a UNet framework to facilitate text-based image generation, establishing itself as a mainstream model in image generation (Podell et al., 2023; Li et al., 2023; Feng et al., 2023). Fine-tuning pre-trained image generation models can enhance their adaptation to specific application scenarios, as seen in techniques like LoRA (Hu et al., 2021) and DreamBooth (Ruiz et al., 2023). For theme control in text-to-image generation, several works (Kumari et al., 2023; Ye et al., 2023; Wang et al., 2024a) focus on custom generation for defined pictorial concepts, with ControlNet (Zhang et al., 2023a) additionally offering control over other modalities such as depth information. These Diffusion models have revolutionized the field of image generation with their powerful generative capabilities. However, the iterative denoising process is completely different from human painting, and cannot be mimicked with a brush, thus offering limited guidance for painting instruction. This paper attempts to use pre-trained diffusion models to generate painting processes that resemble those of artists.
@@ 4.3. Image to Painting Process
In this section, we present the experimental results based on the Artwork Replication Network. As shown in Figure 6(a), by controlling the last frame with a reference image, we have achieved the conversion of artworks into painting processes. In this task, a text prompt is not necessary. We can control different strategies of the painting process by swapping out the painting model or the painting LoRA model, for instance, using oil paint brush strokes to render Van Gogh’s Starry Night or translucent strokes for Chinese ink paintings. When using a semi-finished piece as the reference image to control the first frame, we implemented text-guided painting completion capabilities. As displayed in Figure 6(b), the results from ProcessPainter align well with the text prompts, demonstrating good consistency. Based on the same semi-finished image, different results can be obtained through various text prompts.

### v2

@@ Abstract.
The painting process of artists is inherently stepwise and varies significantly among different painters and styles. Generating detailed, step-by-step painting processes is essential for art education and research, yet remains largely underexplored. Traditional stroke-based rendering methods break down images into sequences of brushstrokes, yet they fall short of replicating the authentic processes of artists, with limitations confined to basic brushstroke modifications. Text-to-image models utilizing diffusion processes generate images through iterative denoising, also diverge substantially from artists’ painting process. To address these challenges, we introduce ProcessPainter, a text-to-video model that is initially pre-trained on synthetic data and subsequently fine-tuned with a select set of artists’ painting sequences using the LoRA model. This approach successfully generates painting processes from text prompts for the first time. Furthermore, we introduce an Artwork Replication Network capable of accepting arbitrary-frame input, which facilitates the controlled generation of painting processes, decomposing images into painting sequences, and completing semi-finished artworks. This paper offers new perspectives and tools for advancing art education and image generation technology. Our code is available at: https://github.com/nicolaus-huang/ProcessPainter.
@@ 2.1. Stroke-based Rendering
The problem of teaching machines “how to paint” has been extensively studied in the context of stroke-based rendering (SBR), which focuses on the recreation of non-photorealistic imagery through appropriate positioning and selection of discrete elements such as paint strokes or stipples(Hertzmann, 2003). Early SBR approaches involved either a greedy, iterative search or required user interaction to address the decomposition problem (Haeberli, 1990; Litwinowicz, 1997; Hertzmann, 2022; Song et al., 2022). In recent years, recurrent neural networks (RNNs) and reinforcement learning (RL) have been extensively employed to generate strokes sequentially (Ha and Eck, 2017; Zhou et al., 2018; Xie et al., 2013; Singh et al., 2022). Moreover, integrating adversarial training (Nakano, 2019) has emerged as a viable method for generating non-deterministic stroke sequences. adversarial updates of actors, critics, and discriminators. Stylized Neural Painting(Zou et al., 2021; Kotovenko et al., 2021) proposed stroke optimization strategy that iteratively searches optimal parameters for each stroke and and is possible to be optimized jointly with neural style transfer. Similarly, the field of vector graphic generation also constructs images using a set of strokes and graphic elements(Frans et al., 2022; Song et al., 2023; Song, 2022; Song and Zhang, 2022) . Although the aforementioned methods achieve the painting process from abstract to specific, these methods still significantly differ from the creative processes of human artists. This is because the painting processes of different artists and subjects vary significantly, making it challenging for a single method to simulate all. Recently, the emergence of Stable Diffusion(Rombach et al., 2022) technology has significantly improved the effectiveness of generative images. Exploring the generation of painting processes based on diffusion models is a promising new field. To this end, we propose ProcessPainter, which fine-tunes diffusion models using data from artists’ painting processes to learn their true distributions.
@@ 2.2. Text2image Diffusion Model
Recent studies have demonstrated that diffusion models are capable of generating high-quality synthetic images, effectively balancing diversity and fidelity. Models based on diffusion models or their variants, such as those documented in (Nichol et al., 2021; Ramesh et al., 2022; Saharia et al., 2022), have successfully addressed the challenges associated with text-conditioned image synthesis. Stable Diffusion (Rombach et al., 2022), a model based on the Latent Diffusion Model, incorporates text conditioning within a UNet framework to facilitate text-based image generation, establishing itself as a mainstream model in image generation (Podell et al., 2023; Li et al., 2023; Feng et al., 2023). Fine-tuning pre-trained image generation models can enhance their adaptation to specific application scenarios, as seen in techniques like LoRA (Hu et al., 2021) and DreamBooth (Ruiz et al., 2023). For theme control in text-to-image generation, several works (Kumari et al., 2023; Ye et al., 2023; Wang et al., 2024a; Zhang et al., 2024a, b) focus on custom generation for defined pictorial concepts, with ControlNet (Zhang et al., 2023a) additionally offering control over other modalities such as depth information. These Diffusion models have revolutionized the field of image generation with their powerful generative capabilities. However, the iterative denoising process is completely different from human painting, and cannot be mimicked with a brush, thus offering limited guidance for painting instruction. This paper attempts to use pre-trained diffusion models to generate painting processes that resemble those of artists.
@@ 4.3. Image to Painting Process
In this section, we present the experimental results based on the Artwork Replication Network. As shown in Fig. 6(a), by controlling the last frame with a reference image, we have achieved the conversion of artworks into painting processes. In this task, a text prompt is not necessary. We can control different strategies of the painting process by swapping out the painting model or the painting LoRA model, for instance, using oil paint brush strokes to render Van Gogh’s Starry Night or translucent strokes for Chinese ink paintings. When using a semi-finished piece as the reference image to control the first frame, we implemented text-guided painting completion capabilities. As displayed in Figure 6(b), the results from ProcessPainter align well with the text prompts, demonstrating good consistency. Based on the same semi-finished image, different results can be obtained through various text prompts.